A sample of magnesium metal reacts completely with water via the following reaction: Mg (s) + 2 H2O (l) → Mg(OH)2 (aq) + H2 (g) The hydrogen produced is collected over water at 25.0°C. If the volume of the gas is 7.80L and the pressure is 0.980 atm, calculate the mass of magnesium metal reacted. (Pwater @ 25.0°C = 23.8 mm Hg)
Answer #1

- 760 mm Hg = 1 catur .., mm Hg = (1760) atro 23.8 mm Hg = (23.8 / 760) atm = 0.0-313-ath - Protal = PH₂O + Pre Lor, Pr = Pro